A man donated a certain amount on a trust fund that will provide an end of the month scholarship grantee for poor deserving students for a period of 10 years. This amount earns an interest rate of 8% compounded monthly. The scholarship amounts to P15,000 every month and the first grant is to be given three years after donation.

A. What was the amount of the actual donation?

B. How much amount of money will be still available on the trust fund, seven years after the beginning of scholarship program?

C. How much total amount was already granted by the program to its scholars for its half-period of operation?
Answer #1

Facts Date of Donation Period of Scholorship Scheme Scholorship Amout Interval of Payment Payment starts from Rate of InteresA What was the amount of the actual donation If the fund pays 15000 per month after three years from the date of donation, thHow much amount of money will be still available on the trust fund, seven years after the beginning of scholarship program? BOpening Balance Interest Total Payment Closing Balance (a) (b=a*15%*1/12) (c=a+b) (e=c-d) Year 4 962,376.00 953,791.84 945,15

Year 5 13 855,503.84 846,207.20 836,848.58 827,427.57 817,943.76 808,396.72 798,786.03 789,111.27 779,372.01 769,567.82 759,6Year 6 25 739,761.35 729,693.09 719,557.71 709,354.76 699,083.80 688,744.35 678,335.98 667,858.22 657,310.61 646,692.68 636,0Year 7 614,412.29 603,508.37 592,531.76 581,481.97 570,358.52 559,160.91 547,888.65 536,541.24 525,118.18 513,618.97 502,043.Year 8 49 478,659.31 466,850.38 454,962.71 442,995.80 430,949.10 418,822.10 406,614.24 394,325.00 381,953.84 369,500.20 356,9Year 9 61 331,638.91 318,849.84 305,975.50 293,015.34 279,968.77 266,835.23 253,614.13 240,304.89 226,906.93 213,419.64 199,8Year 10 172,415.89 158,565.32 144,622.43 130,586.58 116,457.15 102,233.53 87,915.09 73,501.19 58,991.20 44,384.47 29,680.37 1How much total amount was already granted by the program to its scholars for its half-period of operation? Half period of ope